OpenStudy (anonymous):

Convert each subtraction problem to addition and type in the matching number. Answer 3/4 - 2/3 = Read Answer Items for Question 6 3/4 - -2/3 = Read Answer Items for Question 6 -3/4 - 2/3 = Read Answer Items for Question 6 -3/4 - -2/3 = Read Answer Items for Question 6 Answer A. -3/4 + 2/3 B. -3/4 + -2/3 C. 3/4 + -2/3 D. 3/4 + 2/3

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

nope i have no clue

OpenStudy (skullpatrol):

What is the definition of subtraction?

OpenStudy (anonymous):

to take away

OpenStudy (skullpatrol):

Yes, but I'm asking for the algebraic definition.

OpenStudy (anonymous):

subtract   Use Subtract in a sentence sub·tract [suhb-trakt] Show IPA verb (used with object) 1. to withdraw or take away, as a part from a whole. 2. Mathematics . to take (one number or quantity) from another; deduct. verb (used without object) 3. to take away something or a part, as from a whole.

OpenStudy (skullpatrol):

Yes, that is the dictionary definition, in algebra there is a symbolic definition that uses letters instead of numbers.

OpenStudy (anonymous):

can you tell me what it is please

OpenStudy (anonymous):

what the definition is

OpenStudy (skullpatrol):

Let me explain the idea first, to subtract a number b you add the opposite of b.

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

$$\huge a-b=a+(-b)$$

OpenStudy (anonymous):

oh ok that makes more sence thanks

OpenStudy (anonymous):

what if it is a + plus a -

OpenStudy (skullpatrol):

what do you mean?

OpenStudy (anonymous):

sorry a + subtracted from a -

OpenStudy (anonymous):

\[a--b= ?\]

OpenStudy (anonymous):

i will be right back

OpenStudy (skullpatrol):

\[\text{Definition of subtraction: } a-b=a+(-b)\]

OpenStudy (anonymous):

ok i am back

OpenStudy (skullpatrol):